Circular No. 6403

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996aa IN NGC 5557 W. Johnson, Anza, CA, reports his discovery of a supernova (mag ∼

17) on two CCD image exposed on May 16 UT. The new star is located ∼ 5′′ northwest the center of NGC 5557 (α = 14^{h}18^{m}.4, δ = +36^{o}30′, equinox 2000). SN 1996aa is not present on a Carnegie Atlas print or on a slightlyout-of-focus CCD image taken by Johnson in 1995. There is a foreground star of mag ∼ 15 located on the opposite side of the galaxy, ∼ 15′′ southsoutheast from the galaxy nucleus.

P. Garnavich, A. Riess, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram was obtained by E. Barton with the 1.5-m Tillinghast telescope on May 19.26 UT, showing that this is a type-Ia supernova, 6–8 weeks past maximum. SN 1996aa appears 5′′ west and 3′′ north of the galaxy’s center.

COMET C/1995 O1 (HALE-BOPP) D. J. Osip, R. K. Pina, C. M. Telesco, and R. S. Fisher, University

of Florida (UF); and M. S. Hanner, Jet Propulsion Laboratory, report the first mid-infrared detections of comet C/1995 O1, using the UF/NASA imager/spectrometer (OSCIR; field-of-view 28′′ × 28′′) at the Infrared Telescope Facility on May 19.7 UT, with the comet at a heliocentric distance of 4.36 AU. Images were obtained with the broadband N (10.8-μm) and the IHW 18-μm filters. Preliminary data reduction yields flux densities (integrated over a 3′′.25 box aperture) of 1.16 ± 0.02 Jy at 10.8 μm and 3.35 ± 0.19 Jy at 18 μm. The uncertainties represent noise statistics and do not include photometric uncertainties of ∼ 10 percent. These flux densities indicate a bandwidth-corrected color temperature of 224 K, compared to the blackbody equilibrium temperature at 4.36 AU of 133 K.

SUPERNOVA 1996Z IN NGC 2935 S. Nakano, Sumoto, Japan, reports the following precise position for

SN 1996Z measured by Y. Kushida from a CCD image obtained by R. Kushida, Yatsugatake South Base Observatory, on May 18.463 UT: α = 9^{h}36^{m}44^{s}.82, δ = −21^{o}08′51′′.7 (equinox 2000.0). The magnitude of SN 1996Z was measured to be V = 14.2.

1996 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green